OpenStudy (coolshubhs):

If the sum of first n terms of two arithmetic progression are in the ratio (7n + 1) : (4n + 27), find the ratio of their 11th terms.

hartnn (hartnn):

similar to this. http://openstudy.com/study#/updates/50b0a39de4b0e906b4a5fe9d

OpenStudy (mayankdevnani):

very long answer..eh aslo similar http://www.answerbag.com/q_view/1128650

OpenStudy (mayankdevnani):

oh!!!

OpenStudy (mayankdevnani):

ok @coolshubhs

OpenStudy (coolshubhs):

Let the two A.Ps be A.P1 and A.P2 and let the first term and common differenceof A.P1 be a1and d1. and let the first term and common differenceof A.P2 be a2and d2. now. According to question Sum of n terms of A.P1/Sum of n terms of A.P2 = (7n+1)/(4n+27) n/2 [2a1 + (n-1)d1] / n/2 [2a2 + (n-1)d2] = (7n+1)/(4n+27) [2a1 + (n-1)d1] / 2a2 + (n-1)d2] = (7n+1)/(4n+27) ...........(1) Now Ratio of their 11th term. a1 + (11-1)d1/a2 + (11-1)d2 multiply and divide by 2 2a1 + 20d1/2a2 + 20d2..................(2) Comparing (1) and (2) n-1=20 n=21 therefore Ratio of their 11th term. 7(21) + 1 / 4(21) + 27 148/111 Read more: If the sum of first n terms of two arithmetic progression are in the ratio (7n + 1) : (4n + 27), find the ratio of their 11th terms.

OpenStudy (coolshubhs):

but Comparing (1) and (2) ???

OpenStudy (coolshubhs):

??

OpenStudy (coolshubhs):

:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(